OpenStudy (anonymous):

[10.04] What is the exact value of the expression 3the square root of 11. + the square root of 6. - 9the square root of 11.? -5the square root of 17. -6the square root of 11. + the square root of 6. -3the square root of 11. + the square root of 6. -5the square root of 6.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ 3 \sqrt{11} + \sqrt{6}- 9 \sqrt{11} }\]\[\huge\color{blue}{ rig ht? }\]

OpenStudy (anonymous):

Yes!!!!

OpenStudy (solomonzelman):

\[\huge\color{blue}{ 3\sqrt{11}+\sqrt{6}-9 \sqrt{11} }\]\[\huge\color{blue}{ \color{red} { 3\sqrt{11} } +\sqrt{6}\color{red} { -9\sqrt{11} } }\]\[\huge\color{blue}{ \color{red} { 3\sqrt{11} -9 \sqrt{11} } +\sqrt{6} }\]\[\huge\color{blue}{ \color{red} { -6\sqrt{11} } +\sqrt{6} }\]\[\huge\color{blue}{ \color{rblue} { -6\sqrt{11} } +\sqrt{6} }\]

OpenStudy (solomonzelman):

See how I add like terms?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

Good!

OpenStudy (anonymous):

can you help me with this one too [10.04] What is the simplified form of the expression the square root of the quantity 16 times c to the eighth power. - the square root of c squared. + 5the square root of c squared. + the square root of the quantity 4 times c squared.? 8c4 + 8c 4c4 + 8c 8c4 + 6c 4c4 + 6c

OpenStudy (solomonzelman):

Can you draw the question, or write it using equation editor please?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):